Activity continued on the October 1 dome. The banded seismic tremor also continued, although the level of activity has reduced considerably since yesterday. There are still dangers of large pyroclastic flows into the Tar River valley and of sudden collapses from the Galway's Wall. Either could have serious consequences for much of the evacuated area. Depending on the wind direction, rockfalls from the dome may cause some ashfall in areas to the west or northwest of the volcano.
There is still no access to zone A/B. Restricted access is allowed to zone C/D, but only for essential purposes and by people with a means of rapid exit. Zones E, F and G all have normal occupation.
Rockfall and small pyroclastic flow activity from the October 1 dome continued throughout this reporting period. During excellent viewing conditions last night, it was clear that the entire eastern face of the October 1 dome was active. Numerous rockfalls occurred and there was at least one small pyroclastic flow which travelled past the Tar River Soufriere. This generated a moderate-sized ash cloud and ashfall in the Plymouth area. Rockfall and small pyroclastic flow activity continued today with the largest flows again reaching the Tar River Soufriere. Many of these produced small ash clouds which travelled westwards on the wind. Satellite observations at 9:45 am reported ash at a height of 3 - 5,000 ft approximately 60 miles WSW from the volcano.
Seismic activity was again dominated by banded tremor. The peak at about 9pm yesterday was the largest in the sequence to date. Since then the peaks have reduced in size. The interval between the peaks has reduced considerably, to about five hours. Rockfall activity was high in this reporting period, with 44 rockfall events recorded.
Gravity measurements were taken today at stations on the western flank of the volcano. Theodolite measurements were taken to the top of the dome. Initial processing of this data shows few significant changes since yesterday.
EDM measurements to the reflector on Castle Peak were made late yesterday afternoon. The data has yet to be processed.